OpenStudy (anonymous):

Find the antiderivative F of f that satisfies the fiven condition. Check your answer by compairing the graphs of f and F.

OpenStudy (anonymous):

\[f(x)=4-3(1+x^2)^{-1} ; F(1)=0\]

OpenStudy (anonymous):

I thought that by distubuting the 3 into 1+x^2 would make it easier to intergrate

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

ok but thats where i get lost

OpenStudy (helder_edwin):

u have to integrate \[ \large \int(4-3(1+x^2)^{-1})\,dx= \int\left(4-\frac{3}{1+x^2}\right)\,dx \]

OpenStudy (anonymous):

ok can you show me how to intergrate the fraction? I know the 4 once intergrated is 4x

OpenStudy (helder_edwin):

\[ \large =\int4\,dx-3\int\frac{1}{1+x^2}\,dx \] do u know any function whose derivative is 1/(1+x^2) ??

OpenStudy (anonymous):

isn't that tan inverse?

OpenStudy (helder_edwin):

yes!! \[ \large\frac{d}{dx}\arctan x=\frac{1}{1+x^2} \]

OpenStudy (helder_edwin):

then u got it!!!

OpenStudy (anonymous):

ok so just to make sure if i wrote this back in function form it would be \[f(x)=4x-arctanx+c\]

OpenStudy (helder_edwin):

\[ \large F(x)=4x-3\arctan x+c \] yes. now use your initial condition F(1)=0 to find c

OpenStudy (anonymous):

Awesome! Thanks a ton for the help.

OpenStudy (helder_edwin):

u r welcome don't forget to find the value of c